Find the distance between the parallel lines corresponding to y=2x+3 and y=2x+7.

Answer by Alan3354(69443)

Find the distance between the parallel lines corresponding to y=2x+3 and y=2x+7.
---------
m = 2 = tan(angle with x-axis)
4 = hyp of right triangle of the 2 y-ints (0,3) & (0,7)
d = 4*cos(atan(2)) =~ 1.7888 = sqrt(3.2)